Let f(x)= -8x-6, h(x)= x+6/-8
Evaluate (f white circle h)(-62)
The composite function (f white circle h) evaluated at x = -62 is (f white circle h)(-62)=?

Answer by Theo(13342)

f(x) = -8x-6
h(x) = (x+6)/-8
you want to find f(h(x)) when x = -62.
first you find h(-62) = (-62+6)/-8 = -56/-8 = 7
you then find f(h(-62)) = f(7) = -8*7-6 = -56-6 = -62.
your result is f(h(x)) = -62 when x = -62.
in this case, f(x) is the inverse function of (h(x).
note that h(x) results in the coordinate point (-62,7) when x is equal to -62 and that f(x) results in the coordinate point (7,-62) when x = 7.
this is one of the attributes of inverse functions.
here's a reference on inverse functions.
https://mathbitsnotebook.com/Algebra1/Functions/FNInverse.html